# Clustering in Dynamic Environments: A Framework for Benchmark Dataset Generation With Heterogeneous Changes

### 24 Feb 2024 v1

@@ 1. Introduction
Dynamic optimization algorithms (DOAs) have demonstrated their potential in effectively solving various classes of DOPs (yazdani2021DOPsurveyPartA, ). DOAs’ distinct advantage lies in their ability to learn from the past and leverage this historical knowledge, significantly improving and accelerating the optimization process in the current environment. Beyond tracking the moving optimum in DOPs, a subset of DOAs excels in identifying solutions that are robust against imminent environmental changes. These methods, commonly referred to as robust optimization over time (ROOT) (yazdani2023robust, ), further utilize historical knowledge to predict the future quality and robustness of solutions, a necessary feature for DOPs where frequent modifications to the deployed solution are impractical or costly. Such challenges are especially prevalent in dynamic clustering for facility location problems, where frequently changing the locations of facilities (i.e., the cluster centers) is both undesirable and costly.
@@
Meta-heuristic algorithms have established themselves as powerful tools for static clustering tasks, offering flexibility and robustness that are well-suited to finding optimal clustering in a fixed environment (hruschka2009survey, ). Despite the large literature on using meta-heuristic algorithms for clustering in static environments and the fact that many real-world clustering scenarios inherently involve dynamic environments, the application of DOAs in clustering in dynamic environments remains largely unexplored. DOAs extend the capabilities of meta-heuristic algorithms by incorporating historical knowledge, which can be crucial for tackling the dynamic aspects of clustering tasks where data, structures, and even the number of clusters can change over time. The limited application of DOAs in clustering in dynamic environments can be attributed to several factors. A primary challenge is the scarcity of benchmark datasets with the necessary diversity, controllable characteristics, and complexities to effectively test and evolve these algorithms. Unlike the static domain, which benefits from an abundance of datasets, dynamic clustering is yet to develop a comprehensive suite of benchmarks with the required flexibility and variety. Such benchmarks are essential for the progressive advancement of DOAs, enabling them to address a broader spectrum of real-world challenges.
@@
Most existing DOAs are tailored for problems with simplistic assumptions, such as considering only one type of environmental change with uniform temporal and spatial severity and pattern (li2008GDBG, ; grefenstette1999evolvability, ; li2019open, ; branke1999memory, ; yazdani2021DOPsurveyPartB, ; yazdani2020benchmarking, ). This creates a gap between academic research and the complex, multifaceted nature of practical applications, such as clustering in dynamic environments. To bridge this gap and unlock the full potential of DOAs for a wide range of clustering applications in dynamic environments, this paper introduces a new Dynamic Dataset Generator (DDG). This benchmark generator is designed to facilitate rigorous research and aid in the deployment of DOAs for clustering in dynamic environments, paving the way for these algorithms to meet the diverse and complex challenges presented by real-world dynamic settings.
@@
DDG distinguishes itself from previous dynamic benchmark generators (yazdani2021DOPsurveyPartB, ) in the field of DOPs and the dynamic synthetic dataset generators in the field of unsupervised learning (lu2018learning, ) under concept drift by its ability to simulate a wide spectrum of dynamic scenarios. Traditional approaches often simulate simple and mostly homogeneous environmental changes, lacking the complexity observed in real-world dynamics (yazdani2021DOPsurveyPartA, ). In contrast, DDG adeptly simulates a spectrum of changes, from gradual, minor environmental changes to abrupt, significant transformations. This flexibility enables DDG to effectively mimic real-world dynamics, creating scenarios that are heterogeneous in both severity and frequency of changes.
Furthermore, DDG advances beyond the conventional approach of fixed change frequency used in earlier benchmarks. Instead, it treats each change as probabilistic event, occurring with a specific likelihood at each time-tick (function evaluation), thereby offering a more realistic representation of dynamic environments.
@@
In light of these limitations, this paper introduces the Generalized Numerical Benchmark Generator (GNBG), a configurable, flexible, and user-friendly tool explicitly designed to embody the desirable properties of an effective benchmark suite. Unlike conventional benchmark suites that rely on a variety of fixed mathematical functions, GNBG employs a single, parametric baseline function capable of generating a diverse range of problem instances with controllable characteristics and levels of difficulty. By manipulating various parameters within GNBG, users gain the ability to tailor the properties of the generated problem instances, including:
DDG is a configurable benchmark generator, offering users the flexibility to customize problem instances by adjusting various parameters within the generator. It exhibits several key properties essential for an effective benchmark, as outlined in (bartz2020benchmarking, ). These properties include the capability to generate a diverse range of problem instances with controllable dynamic characteristics and distributions (diversity) and controllable challenges and levels of difficulty (complexity variety). Furthermore, DDG simulates dynamic scenarios pertinent to various types of clustering in real-world dynamic environments, such as data with concept drift and facility location problems (relevance and representativity). Additionally, DDG is scalable in terms of the number of variables (scalability), and its source code is made publicly available at (yazdani2024DDGcode, ) (accessability). Note, however, that one desirable property not encompassed by DDG is the availability of known optimal solutions. This is primarily due to the inherent nature of clustering problems as a domain of unsupervised learning: the absence of a known optimum solution or definitive ground truth for determining optimal clustering (hruschka2009survey, ). This necessitates a comparative analysis approach, focusing on the relative performance of algorithms when using DDG for benchmarking algorithms.
@@ 2. Background and Related Work
To address the limitations in existing benchmark problems that hinder the progression of DOAs, a direct leap into solving real-world problems with their multifaceted challenges and indeterminate characteristics is not feasible. Instead, there is a pressing need for advanced benchmarks that offer more realistic landscapes and dynamics. Clustering in dynamic environments emerges as an ideal candidate for this purpose. It represents a subset of DOPs where clustering solutions must be continually updated to adapt to environmental changes. These problems cover a wide range of challenges and applications, from adapting to data exhibiting concept drift to optimizing real-time resource allocation in facility location problems. They call for sophisticated approaches capable of responding efficiently to ongoing changes, ensuring that clustering solutions stay relevant and effective in a dynamic setting.
@@
In the field of clustering in static environments, there is an abundance of datasets available for evaluating algorithms. However, the landscape is markedly different for clustering in dynamic environments. Most benchmark datasets designed for learning under concept drift predominantly cater to classification problems, where the labels may vary over time (gama2014survey, ). Moreover, these datasets often fail to encompass a broad spectrum of dynamic scenarios. Firstly, the configurability of many synthetic datasets is limited. Parameters such as the number of variables and classes/clusters are typically fixed in many cases (lu2018learning, ), though there are a few instances where they offer the flexibility to control the number of variables and classes/clusters (bifet2010moa, ). More significantly, the dynamics represented in these datasets tend to be restricted, focusing more on dataset-specific characteristics rather than systematic and controllable changes in underlying distributions.
@@ 3. Proposed Dynamic Dataset Generator
This section introduces the Dynamic Dataset Generator (DDG) [^1] The source code of DDG is publicly available and can be accessed at (yazdani2024DDGcode, )., a novel tool tailored for the generation of dynamic datasets with controllable characteristics. Subsequent subsections provide detailed descriptions of DDG’s key components: the data generator and dynamics.
@@ 3.1. Data Generator
(1) $\mathcal{N}_{i}^{(t)}\left(\mathbf{c}_{i}^{(t)},\bm{\sigma}_{i}^{(t)},\bm{\Theta}^{(t)}\right)=\mathbf{r}\bm{\sigma}_{i}^{(t)}\mathbb{R}\left(\bm{\Theta}^{(t)}\right)+\mathbf{c}_{i}^{(t)},$
where $\mathcal{N}_{i}^{(t)}\left(\mathbf{c}_{i}^{(t)},\bm{\sigma}_{i}^{(t)},\bm{\Theta}^{(t)}\right)$ represents the $i$th DGC at time $t$. In this equation, $\mathbf{r}$ is a $d^{(t)}$-dimensional random vector, with each element drawn from a standard normal distribution $\mathcal{N}(0,1)$. The vector $\mathbf{c}_{i}^{(t)}$ denotes the center position (i.e., mean) of the $i$th DGC in the $t$th environment, and $d^{(t)}$ represents the number of variables at time $t$. The vector $\bm{\sigma}_{i}^{(t)}$ is a $d^{(t)}$-dimensional vector representing the standard deviation, or the spread of the distribution, for the $i$th DGC across each dimension. The matrix $\bm{\Theta}_{i}^{(t)}$, a $d^{(t)}\times d^{(t)}$ matrix, plays a crucial role in defining the rotation of each component. All elements on and below the principal diagonal of $\bm{\Theta}_{i}^{(t)}$ are set to zero. Each off-diagonal element above the diagonal in the $j$th row and $k$th column (where $j<k$) specifies a rotation angle. This angle determines the degree of rotation for the projection of every point generated by the $i$th DGC in environment $t$, specifically onto the $x_{j}-x_{k}$ plane.
@@ 21 Input: $d^{(t)}$ and $\bm{\Theta}_{i}^{(t)}$
(a) Probability contour plot of a DDG with three DGCs.
@@ 3.2.1. Gradual Local Changes
Figure 4 illustrates the behavior of the dynamics presented in Equations (5) to (7), particularly how they respond to variations in the probability of altering the direction factor. This probability factor serves a role analogous to $\rho$ in Equation (4), as both parameters influence the correlation of changes over time. Note that when the correlation is minimal (representing a fully random change) and the change severity is low, the parameters tend to fluctuate slightly around their initial values without significant cumulative change, as shown in Figure 4(d). This observation underscores the importance of employing more correlated changes in parameters to simulate scenarios with frequent yet gradual changes, a characteristic common in many real-world problems.
@@ 4. Discussion: Performance Measures, Settings, and Preliminary Findings
The DDG framework has been meticulously developed to benchmark algorithms designed for clustering in dynamic environments. It encompasses a comprehensive array of configurable parameters, facilitating the simulation of diverse dynamic scenarios tailored to specific research objectives. Table 1 in the appendix summarizes all parameters of DDG, and provides suggested values and ranges for these parameters.
@@ 5. Conclusion
In this paper, we introduced the Dynamic Dataset Generator (DDG), a comprehensive benchmark generator tool crafted for creating dynamic datasets with known and controllable characteristics, specifically tailored for the evaluation of algorithms in clustering in dynamic environments. At the heart of DDG are multiple Dynamic Gaussian Components (DGCs), which facilitate a broad spectrum of dynamic scenarios. These range from subtle, continuous environmental changes to abrupt, substantial transformations. DDG is uniquely equipped to generate dynamic scenarios characterized by heterogeneous changes. These changes encompass a variety of temporal and spatial severities, patterns, and areas of influence, effectively capturing both local and global dynamics. Through this diverse and flexible approach, DDG serves as a vital tool for realistically simulating and assessing the performance of clustering algorithms in dynamic, real-world-like scenarios.
@@
In the appendix, we delve into the intricacies of the Dynamic Dataset Generator (DDG), a tool that stands as a discrete-event-based simulation capable of replicating a diverse range of dynamic scenarios. The core of DDG’s versatility lies in its capacity to simulate various dynamics, each governed by a distinct set of parameters. This multitude of parameters, while seemingly complex, is indispensable for the accurate representation and simulation of a wide range of scenarios. The complexity of real-world dynamics and the need for their precise emulation in a controlled environment necessitate this extensive parameterization. By allowing fine-grained control over each aspect of the simulation, DDG provides researchers with the flexibility to explore, test, and validate a wide spectrum of scenarios, each with its unique characteristics and challenges. Table 1 summarizes all the parameters used in DDG along with their suggested ranges and values. To disable any type of dynamic in DDG, researchers can set the likelihood parameter ($\mathfrak{p}$) of that dynamic to zero. Additionally, to avoid changing a specific parameter, its change severity value can be set to zero.

### 09 Apr 2024 v2

@@ 1. Introduction
Dynamic optimization algorithms (DOAs) have demonstrated their potential in effectively solving various classes of DOPs (yazdani2021DOPsurveyPartA, ; yazdani2018thesis, ). DOAs’ distinct advantage lies in their ability to learn from the past and leverage this historical knowledge, significantly improving and accelerating the optimization process in the current environment. Beyond tracking the moving optimum in DOPs, a subset of DOAs excels in identifying solutions that are robust against imminent environmental changes. These methods, commonly referred to as robust optimization over time (ROOT) (yazdani2023robust, ), further utilize historical knowledge to predict the future quality and robustness of solutions, a necessary feature for DOPs where frequent modifications to the deployed solution are impractical or costly. Such challenges are especially prevalent in dynamic clustering for facility location problems, where frequently changing the locations of facilities (i.e., the cluster centers) is both undesirable and costly.
@@
Meta-heuristic algorithms have established themselves as powerful tools for static clustering tasks, offering flexibility and robustness that are well-suited to finding optimal clustering in a fixed environment (hruschka2009survey, ). Despite the large literature on using meta-heuristic algorithms for clustering in static environments and the fact that many real-world clustering scenarios inherently involve dynamic environments, the application of DOAs in clustering in dynamic environments remains largely unexplored. DOAs extend the capabilities of meta-heuristic algorithms by incorporating historical knowledge, which can be crucial for tackling the dynamic aspects of clustering tasks where data, structures, and even the number of clusters can change over time. The limited application of DOAs in clustering in dynamic environments can be attributed to several factors. A primary challenge is the lack of benchmark datasets with the necessary diversity, controllable characteristics, and complexities to effectively test and evolve these algorithms. Unlike the static domain, which benefits from an abundance of datasets, dynamic clustering is yet to develop a comprehensive suite of benchmarks with the required flexibility and variety. Such benchmarks are essential for the progressive advancement of DOAs, enabling them to address a broader range of real-world challenges.
@@
Most existing DOAs are tailored for problems with simplistic assumptions, such as considering only one type of environmental change with uniform temporal and spatial severity and pattern (li2008GDBG, ; grefenstette1999evolvability, ; li2019open, ; branke1999memory, ; yazdani2021DOPsurveyPartB, ; yazdani2020benchmarking, ). This creates a gap between academic research and the complex nature of practical applications, such as clustering in dynamic environments. To bridge this gap and unlock the full potential of DOAs for a wide range of clustering applications in dynamic environments, this paper introduces a new Dynamic Dataset Generator (DDG). This benchmark generator is designed to facilitate rigorous research and aid in the deployment of DOAs for clustering in dynamic environments, paving the way for these algorithms to meet the diverse and complex challenges presented by real-world dynamic settings.
@@
DDG distinguishes itself from previous dynamic benchmark generators (yazdani2021DOPsurveyPartB, ) in the field of DOPs and the dynamic synthetic dataset generators in the field of unsupervised learning (lu2018learning, ) under concept drift by its ability to simulate a wide range of dynamic scenarios. Traditional approaches often simulate simple and mostly homogeneous environmental changes, lacking the complexity observed in real-world dynamics (yazdani2021DOPsurveyPartA, ). In contrast, DDG adeptly simulates a range of changes, from gradual, minor environmental changes to abrupt, significant transformations. This flexibility enables DDG to effectively mimic real-world dynamics, creating scenarios that are heterogeneous in both severity and frequency of changes.
Furthermore, DDG advances beyond the conventional approach of fixed change frequency used in earlier benchmarks. Instead, it treats each change as probabilistic event, occurring with a specific likelihood at each time-tick (function evaluation), thus offering a more realistic representation of dynamic environments.
@@
DDG is a configurable benchmark generator, offering users the flexibility to customize problem instances by adjusting various parameters within the generator. It exhibits several key properties essential for an effective benchmark, as outlined in (bartz2020benchmarking, ). These properties include the capability to generate a diverse range of problem instances with controllable dynamic characteristics and distributions (diversity) and controllable challenges and levels of difficulty (complexity variety). Furthermore, DDG simulates dynamic scenarios pertinent to various types of clustering in real-world dynamic environments, such as data with concept drift and facility location problems (relevance and representativity). Additionally, DDG is scalable in terms of the number of variables (scalability), and its source code is made publicly available at (yazdani2024DDGcodeMATLAB, ; yazdani2024DDGcodePython, ) (accessability). Note, however, that one desirable property not encompassed by DDG is the availability of known optimal solutions. This is primarily due to the inherent nature of clustering problems as a domain of unsupervised learning: the absence of a known optimum solution or definitive ground truth for determining optimal clustering (hruschka2009survey, ). This necessitates a comparative analysis approach, focusing on the relative performance of algorithms when using DDG for benchmarking algorithms.
@@ 2. Background and Related Work
To address the limitations in existing benchmark problems that hinder the progression of DOAs, a direct transition into solving real-world problems with their complex challenges and indeterminate characteristics is not feasible. Instead, there is a pressing need for advanced benchmarks that offer more realistic landscapes and dynamics. Clustering in dynamic environments emerges as an ideal candidate for this purpose. It represents a subset of DOPs where clustering solutions must be continually updated to adapt to environmental changes. These problems cover a wide range of challenges and applications, from adapting to data exhibiting concept drift to optimizing real-time resource allocation in facility location problems. They call for sophisticated approaches capable of responding efficiently to ongoing changes, ensuring that clustering solutions stay relevant and effective in a dynamic setting.
@@
In the field of clustering in static environments, there is an abundance of datasets available for evaluating algorithms. However, the landscape is markedly different for clustering in dynamic environments. Most benchmark datasets designed for learning under concept drift predominantly cater to classification problems, where the labels may vary over time (gama2014survey, ). Moreover, these datasets often fail to encompass a broad range of dynamic scenarios. Firstly, the configurability of many synthetic datasets is limited. Parameters such as the number of variables and classes/clusters are typically fixed in many cases (lu2018learning, ), though there are a few instances where they offer the flexibility to control the number of variables and classes/clusters (bifet2010moa, ). More significantly, the dynamics represented in these datasets tend to be restricted, focusing more on dataset-specific characteristics rather than systematic and controllable changes in underlying distributions.
@@ 3. Proposed Dynamic Dataset Generator
This section introduces the Dynamic Dataset Generator (DDG) [^1] The Python and MATLAB source codes of DDG are publicly available and can be accessed at (yazdani2024DDGcodePython, ) and (yazdani2024DDGcodeMATLAB, ), respectively., a novel tool tailored for the generation of dynamic datasets with controllable characteristics. Subsequent subsections provide detailed descriptions of DDG’s key components: the data generator and dynamics.
@@ 3.1. Data Generator
(1) $\mathcal{N}_{i}^{(t)}\left(\mathbf{c}_{i}^{(t)},\bm{\sigma}_{i}^{(t)},\bm{\Theta}^{(t)}_{i}\right)=\mathbf{r}\bm{\sigma}_{i}^{(t)}\mathbb{R}\left(\bm{\Theta}^{(t)}_{i}\right)+\mathbf{c}_{i}^{(t)},$
where $\mathcal{N}_{i}^{(t)}\left(\mathbf{c}_{i}^{(t)},\bm{\sigma}_{i}^{(t)},\bm{\Theta}^{(t)}_{i}\right)$ represents the $i$th DGC at time $t$. In this equation, $\mathbf{r}$ is a $d^{(t)}$-dimensional random vector, with each element drawn from a standard normal distribution $\mathcal{N}(0,1)$. The vector $\mathbf{c}_{i}^{(t)}$ denotes the center position (i.e., mean) of the $i$th DGC in the $t$th environment, and $d^{(t)}$ represents the number of variables at time $t$. The vector $\bm{\sigma}_{i}^{(t)}$ is a $d^{(t)}$-dimensional vector representing the standard deviation, or the spread of the distribution, for the $i$th DGC across each dimension. The matrix $\bm{\Theta}_{i}^{(t)}$, a $d^{(t)}\times d^{(t)}$ matrix, plays a crucial role in defining the rotation of each component. All elements on and below the principal diagonal of $\bm{\Theta}_{i}^{(t)}$ are set to zero. Each off-diagonal element above the diagonal in the $j$th row and $k$th column (where $j<k$) specifies a rotation angle. This angle determines the degree of rotation for the projection of every point generated by the $i$th DGC in environment $t$, specifically onto the $x_{j}-x_{k}$ plane.
@@ 21 Input: $d^{(t)}$ and $\bm{\Theta}_{i}^{(t)}$
(a) Probability contour plot of a DDG with three DGCs. The warmer the color, the higher the probability; colder colors indicate lower probabilities.
@@ 3.2.1. Gradual Local Changes
Figure 4 illustrates the behavior of the dynamics presented in Equations (5) to (7), particularly how they respond to variations in the probability of altering the direction factor. This probability factor serves a role analogous to $\rho$ in Equation (4), as both parameters influence the correlation of changes over time. Note that when the correlation is minimal (representing a fully random change) and the change severity is low, the parameters tend to fluctuate slightly around their initial values without significant cumulative change, as shown in Figure 4(d). This observation highlights the importance of employing more correlated changes in parameters to simulate scenarios with frequent yet gradual changes, a characteristic common in many real-world problems.
@@ 4. Discussion: Performance Measures, Settings, and Preliminary Findings
The DDG framework has been carefully developed to benchmark algorithms designed for clustering in dynamic environments. It encompasses a comprehensive array of configurable parameters, facilitating the simulation of diverse dynamic scenarios tailored to specific research objectives. Table 1 in the appendix summarizes all parameters of DDG, and provides suggested values and ranges for these parameters.
@@ 5. Conclusion
In this paper, we introduced the Dynamic Dataset Generator (DDG), a comprehensive benchmark generator tool crafted for creating dynamic datasets with known and controllable characteristics, specifically tailored for the evaluation of algorithms in clustering in dynamic environments. At the heart of DDG are multiple Dynamic Gaussian Components (DGCs), which facilitate a broad range of dynamic scenarios. These range from subtle, continuous environmental changes to abrupt, substantial transformations. DDG is uniquely equipped to generate dynamic scenarios characterized by heterogeneous changes. These changes encompass a variety of temporal and spatial severities, patterns, and areas of influence, effectively capturing both local and global dynamics. Through this diverse and flexible approach, DDG serves as a vital tool for realistically simulating and assessing the performance of clustering algorithms in dynamic, real-world-like scenarios.
@@
In the appendix, we explore the parameter setting of the Dynamic Dataset Generator (DDG), a tool that stands as a discrete-event-based simulation capable of replicating a diverse range of dynamic scenarios. The core of DDG’s versatility lies in its capacity to simulate various dynamics, each governed by a distinct set of parameters. This multitude of parameters, while seemingly complex, is indispensable for the accurate representation and simulation of a wide range of scenarios. The complexity of real-world dynamics and the need for their precise emulation in a controlled environment necessitate this extensive parameterization. By allowing fine-grained control over each aspect of the simulation, DDG provides researchers with the flexibility to explore, test, and validate a wide range of scenarios, each with its unique characteristics and challenges. Table 1 summarizes all the parameters used in DDG along with their suggested ranges and values. To disable any type of dynamic in DDG, researchers can set the likelihood parameter ($\mathfrak{p}$) of that dynamic to zero. Additionally, to avoid changing a specific parameter, its change severity value can be set to zero.